# Sharp non-uniqueness for the Navier–Stokes equations in scaling critical spaces

### v1

@@ 4.1. Notations and key lemmas
$F_{2;j_{1},j_{2}}:={}V^{({\rm p})}_{j_{1}}\otimes V^{({\rm p})}_{j_{2}}$ (4.123)
$\quad=2\lambda_{j_{1}-1}^{3}\lambda_{j_{2}-1}^{3}\sum_{k_{1},k_{2}\in\mathcal{K}}\Gamma_{j_{1},k_{1}}[V_{j_{1}-1}]\Gamma_{j_{2},k_{2}}[V_{j_{2}-1}]\Psi_{j_{1},k_{1}}\Psi_{j_{2},k_{2}}k_{1}\otimes k_{2}$ (4.124)
$\quad=\frac{\lambda_{j_{1}-1}^{3}\lambda_{j_{2}-1}^{3}}{\sqrt{h_{j_{1}}h_{j_{2}}}}\sum_{\pm}\sum_{\ell_{1}\in\Lambda_{j_{1}}}\sum_{\ell_{2}\in\Lambda_{j_{2}}}\sum_{k_{1},k_{2}\in\mathcal{K}}\frac{1}{\sqrt{\ell_{1}\ell_{2}}}\Gamma_{j_{1},k_{1}}[V_{j_{1}-1}](x)\Gamma_{j_{2},k_{2}}[V_{j_{2}-1}](x)$ (4.125)
@@
As $F_{2;j_{1},j_{2}}=F_{2;j_{2},j_{1}}$, it suffices to consider the case $j_{1}<j_{2}$. Since it holds
$|\lambda^{\ell_{1}-\ell_{2}}a_{k_{1}}\pm a_{k_{2}}|\geqslant|a_{k_{2}}|-\lambda^{\ell_{1}-\ell_{2}}|a_{k_{1}}|\geqslant\frac{1}{2}|a_{k_{2}}|>0,$ (4.127)
@@
${\left\|F_{2;j_{1},j_{2}}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}\leqslant{}$ $C\lambda_{j_{1}-1}^{3-2\mu}\lambda_{j_{2}-1}^{3-2\mu}\sum_{\ell_{1}\in\Lambda_{j_{1}}}\sum_{\ell_{2}\in\Lambda_{j_{2}}}\sum_{k_{1},k_{2}\in\mathcal{K}}$ (4.128)
@@
$\left.\quad+\frac{\lambda^{-2\ell_{2}}}{\sqrt{\ell_{1}\ell_{2}}}|\lambda^{\ell_{1}-\ell_{2}}a_{k_{1}}\pm a_{k_{2}}|^{-2}{\left\|\nabla^{n}(\Gamma_{j_{1},k_{1}}[V_{j_{1}-1}]\Gamma_{j_{2},k_{2}}[V_{j_{2}-1}]\psi_{j_{1}})\right\|}_{L^{r}}\right)$ (4.130)
$\leqslant{}$ $C\lambda^{\frac{n}{r}j_{1}}\lambda_{j_{1}-1}^{3-2\mu}\lambda_{j_{2}-1}^{3-2\mu}\sum_{\ell_{1}\in\Lambda_{j_{1}}}\sum_{\ell_{2}:\ell_{2}\geqslant\ell_{1}}\lambda^{(\frac{n}{r}-2)\ell_{2}}$ (4.131)
$+C\lambda^{nj_{1}}\lambda_{j_{1}-1}^{3-2\mu}\lambda_{j_{2}-1}^{3-2\mu+2n}\sum_{\ell_{1}\in\Lambda_{j_{1}}}\sum_{\ell_{2}:\ell_{2}\geqslant\ell_{1}}\lambda^{-2\ell_{2}}$ (4.132)
$\leqslant{}$ $C\lambda^{\frac{n}{r}j_{1}}\lambda_{j_{1}-1}^{3-(4-\frac{n}{r})\mu}\lambda_{j_{2}-1}^{3-2\mu}+C\lambda_{j_{1}-1}^{3-4\mu}\lambda_{j_{2}-1}^{3-2\mu+2n}$ (4.133)
$\leqslant{}$ $C\lambda^{\frac{n}{r}j_{1}}\lambda_{j_{1}-1}^{3-(4-\frac{n}{r})\mu}\lambda_{j_{2}-1}^{3-2\mu}.$ (4.134)
@@
$\leqslant C\sum_{j_{1}=1}^{\infty}\lambda^{\frac{n}{r}j_{1}}\lambda_{j_{1}-1}^{3-(4-\frac{n}{r})\mu}\sum_{j_{2}=j_{1}+1}^{\infty}\lambda_{j_{2}-1}^{3-2\mu}$ (4.136)
$\leqslant C\sum_{j_{1}=1}^{\infty}\lambda_{j_{1}-1}^{6-(6-\frac{n}{r})\mu}$ (4.137)
$\leqslant C\lambda^{6-(6-\frac{n}{r})\mu}\leqslant C\lambda^{-\beta}.$ (4.138)
@@
$\leqslant C\left(\lambda^{3-(3-\frac{n}{p_{*}})\mu}+\lambda_{1}^{\frac{n}{r}-2}\lambda^{3-2\mu}\right)\lambda_{1}^{\frac{n}{r}-2}\lambda^{3-2\mu}$ (4.140)
$\leqslant C\lambda_{1}^{\frac{n}{r}-2}\lambda^{6-4\mu}\leqslant C\lambda^{-\beta}.$ (4.141)
@@
We separate the proof into two steps; the first step achieves the construction of the non-trivial solutions to the unforced stationary Navier–Stokes equations, and the second step establishes the desired non-stationary flow by the standard stability argument around the steady state provided in the previous step.
Step 1. Steady Navier–Stokes flow around $V_{\lambda}$. Fix a $n<p_{*}<2n$ and $n/2<r<n$ so that (2.7) with $p=p_{*}$. Let us construct the solution to (1.2) of the form $U_{\lambda}=V_{\lambda}+W_{\lambda}$. To this end, we consider the stationary perturbed system
@@
and define an operator $\mathcal{X}$ on $\dot{B}_{r,1}^{n/r-1}$ by
$\mathcal{X}[W]:=-(-\Delta)^{-1}F-(-\Delta)^{-1}\mathbb{P}\operatorname{div}\left(V_{\lambda}\otimes W+W\otimes V_{\lambda}+W\otimes W\right).$ (5.2)
@@
$\begin{aligned} {\left\|\mathcal{X}[W]\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}&\leqslant C{\left\|F\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-3}}+C\left({\left\|V_{\lambda}\right\|}_{\dot{B}_{p_{*},1}^{\frac{n}{p_{*}}-1}}+{\left\|W\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\right){\left\|W\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}},\\ &\leqslant C_{1}\lambda^{-\beta}+C_{1}\left(\lambda^{-\alpha(p_{*})}+{\left\|W\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\right){\left\|W\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}},\end{aligned}$ (5.3)
@@
${\left\|\mathcal{X}[W]\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\leqslant C_{1}\lambda^{-\beta},\qquad{\left\|\mathcal{X}[W_{1}]-\mathcal{X}[W_{2}]\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\leqslant\frac{1}{2}{\left\|W_{1}-W_{2}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}$ (5.8)
@@
To construct the perturbed solution, we define a map
@@
$\quad\leqslant{}C{\left\|w_{0}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}+C\left({\left\|w_{0}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}+{\left\|w\right\|}_{\widetilde{L^{\infty}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1})}\right){\left\|w\right\|}_{\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}})}$ (5.14)
$\quad\leqslant{}C_{2}{\left\|U_{\lambda}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}+C_{2}\left(\lambda^{-\alpha(p)}+{\left\|w\right\|}_{\widetilde{L^{\infty}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1})}\right){\left\|w\right\|}_{\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}})}$ (5.15)
@@
Finally, we shall prove the asymptotic behavior $u_{\lambda}(t)-U_{\lambda}=w_{\lambda}(t)\to 0$ strongly in $\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ as $t\to\infty$. To see this, we fix a $T>0$ and rewrite the integral equation for $w_{\lambda}$ as

### v2

@@ 4.1. Notations and key lemmas
$F_{2;j_{1},j_{2}}(x):={}V^{({\rm p})}_{j_{1}}\otimes V^{({\rm p})}_{j_{2}}(x)$ (4.127)
$\quad=2\lambda_{j_{1}-1}^{4}\lambda_{j_{2}-1}^{4}\sum_{k_{1},k_{2}\in\mathcal{K}}\sum_{\ell_{1},\ell_{2}\in\Lambda_{j}}\Phi_{j_{1},k_{1},\ell_{1}}(x)\otimes\Phi_{j_{2},k_{2},\ell_{2}}(x)$ (4.128)
$\quad=\frac{\lambda_{j_{1}-1}^{4}\lambda_{j_{2}-1}^{4}}{\sqrt{h_{j_{1}}h_{j_{2}}}}\sum_{\pm}\sum_{\ell_{1}\in\Lambda_{j_{1}}}\sum_{\ell_{2}\in\Lambda_{j_{2}}}\sum_{k_{1},k_{2}\in\mathcal{K}}\frac{1}{\sqrt{\ell_{1}\ell_{2}}}\Gamma_{j_{1},k_{1}}[V_{j_{1}-1}](x)\Gamma_{j_{2},k_{2}}[V_{j_{2}-1}](x)$ (4.129)
@@
As $F_{2;j_{1},j_{2}}=(F_{2;j_{2},j_{1}})^{\top}$, it suffices to consider the case $j_{1}<j_{2}$. Since it holds
$|\lambda^{\ell_{1}-\ell_{2}}a_{k_{1}}\pm a_{k_{2}}|\geqslant 1-\lambda^{\ell_{1}-\ell_{2}}\geqslant\frac{1}{2}>0,$ (4.131)
@@
${\left\|F_{2;j_{1},j_{2}}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}\leqslant{}$ $C\lambda_{j_{1}-1}^{4-2\mu}\lambda_{j_{2}-1}^{4-2\mu}\sum_{\ell_{1}\in\Lambda_{j_{1}}}\sum_{\ell_{2}\in\Lambda_{j_{2}}}\sum_{k_{1},k_{2}\in\mathcal{K}}$ (4.132)
@@
$\left.\quad+\frac{\lambda^{-2\ell_{2}}}{\sqrt{\ell_{1}\ell_{2}}}|\lambda^{\ell_{1}-\ell_{2}}a_{k_{1}}\pm a_{k_{2}}|^{-2}{\left\|\nabla^{n}(\Gamma_{j_{1},k_{1}}[V_{j_{1}-1}]\Gamma_{j_{2},k_{2}}[V_{j_{2}-1}]\psi_{j_{1}})\right\|}_{L^{r}\cap L^{1}}\right)$ (4.134)
$\leqslant{}$ $C\lambda^{\frac{n}{r}j_{1}}\lambda_{j_{1}-1}^{4-2\mu}\lambda_{j_{2}-1}^{4-2\mu}\sum_{\ell_{1}\in\Lambda_{j_{1}}}\sum_{\ell_{2}:\ell_{2}\geqslant\ell_{1}}\lambda^{(\frac{n}{r}-2)\ell_{2}}$ (4.135)
$+C\lambda^{nj_{1}}\lambda_{j_{1}-1}^{4-2\mu}\lambda_{j_{2}-1}^{4-2\mu+2n}\sum_{\ell_{1}\in\Lambda_{j_{1}}}\sum_{\ell_{2}:\ell_{2}\geqslant\ell_{1}}\lambda^{-2\ell_{2}}$ (4.136)
$\leqslant{}$ $C\lambda^{\frac{n}{r}j_{1}}\lambda_{j_{1}-1}^{4-(4-\frac{n}{r})\mu}\lambda_{j_{2}-1}^{4-2\mu}+C\lambda^{nj_{1}}\lambda_{j_{1}-1}^{4-4\mu}\lambda_{j_{2}-1}^{4-2\mu+2n}$ (4.137)
$\leqslant{}$ $C\lambda^{nj_{1}}\lambda_{j_{1}-1}^{4-(4-\frac{n}{r})\mu}\lambda_{j_{2}-1}^{4-2\mu}.$ (4.138)
@@
$\leqslant C\sum_{j_{1}=1}^{\infty}\lambda^{nj_{1}}\lambda_{j_{1}-1}^{4-(4-\frac{n}{r})\mu}\sum_{j_{2}=j_{1}+1}^{\infty}\lambda_{j_{2}-1}^{4-2\mu}$ (4.140)
$\leqslant C\sum_{j_{1}=1}^{\infty}\lambda^{nj_{1}}\lambda_{j_{1}-1}^{8-(6-\frac{n}{r})\mu}$ (4.141)
$\leqslant C\lambda^{8+n-(6-\frac{n}{r})\mu}\leqslant C\lambda^{4n+8-4\mu}=C\lambda^{-\beta}.$ (4.142)
@@
$\leqslant C\left(\lambda^{\frac{n}{p_{*}}+4-(3-\frac{n}{p_{*}})\mu}+\lambda^{n+6-(4-\frac{n}{r})\mu}\right)\lambda^{n+6-(4-\frac{n}{r})\mu}$ (4.144)
$\leqslant C\left(\lambda^{\frac{9}{2}-\frac{5}{2}\mu}+\lambda^{n+6-2\mu}\right)\lambda^{n+6-2\mu}$ (4.145)
$\leqslant C\lambda^{2n+12-4\mu}<C\lambda^{4n+8-4\mu}=C\lambda^{-\beta}.$ (4.146)
@@
We separate the proof into two steps; the first step achieves the construction of the non-trivial solutions to the unforced stationary Navier–Stokes equations, and the second step establishes the desired non-stationary flow by the standard stability argument around the steady state provided in the previous step. Let $V_{\lambda}$ and $F_{\lambda}$ be vector fields constructed in Theorem 4.1.
Step 1. Steady Navier–Stokes flow around $V_{\lambda}$. Fix a $n<p_{*}<2n$ and $n/2<r<n$ so that (2.7) holds with $p$ replaced by $p_{*}$. Let us construct the solution to (1.2) of the form $U_{\lambda}=V_{\lambda}+W_{\lambda}$. To this end, we consider the stationary perturbed system
@@
and define an operator $\mathcal{X}$ on $\dot{B}_{r,1}^{n/r-1}(\mathbb{R}^{n})$ by
$\mathcal{X}[W]:=-(-\Delta)^{-1}F_{\lambda}-(-\Delta)^{-1}\mathbb{P}\operatorname{div}\left(V_{\lambda}\otimes W+W\otimes V_{\lambda}+W\otimes W\right).$ (5.2)
@@
$\begin{aligned} {\left\|\mathcal{X}[W]\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}&\leqslant C{\left\|F_{\lambda}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-3}}+C\left({\left\|V_{\lambda}\right\|}_{\dot{B}_{p_{*},1}^{\frac{n}{p_{*}}-1}}+{\left\|W\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\right){\left\|W\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}},\\ &\leqslant C_{1}\lambda^{-\beta}+C_{1}\left(\lambda^{-\alpha(p_{*})}+{\left\|W\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\right){\left\|W\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}},\end{aligned}$ (5.3)
@@
${\left\|\mathcal{X}[W]\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\leqslant 2C_{1}\lambda^{-\beta},\qquad{\left\|\mathcal{X}[W_{1}]-\mathcal{X}[W_{2}]\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\leqslant\frac{1}{2}{\left\|W_{1}-W_{2}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}$ (5.8)
@@
To construct the perturbed solution, we define a mapping
@@
$\quad\leqslant{}C{\left\|w_{0}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}+C\left({\left\|U_{\lambda}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}+{\left\|w\right\|}_{\widetilde{L^{\infty}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1})}\right){\left\|w\right\|}_{\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}})}$ (5.14)
$\quad\leqslant{}C_{2}{\left\|w_{0}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}+C_{2}\left(\lambda^{-\alpha(p)}+{\left\|w\right\|}_{\widetilde{L^{\infty}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1})}\right){\left\|w\right\|}_{\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}})}$ (5.15)
@@
Finally, we shall prove the asymptotic behavior $u_{\lambda}(t)-U_{\lambda}=w_{\lambda}(t)\to 0$ strongly in $\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ as $t\to\infty$. Let us fix a $T>0$ and rewrite the integral equation for $w_{\lambda}$ as